OpenStudy (anonymous):

Evaluate the expression. 48 ÷ 4 · 6

OpenStudy (texaschic101):

division then multiplication

OpenStudy (texaschic101):

the reason I say this is because division and multiplication are on the same level so you just start from the left and go to the right

OpenStudy (texaschic101):

48/4 * 6 = 12 * 6 = 72

OpenStudy (anonymous):

Yepp^